OpenStudy (alexatiger):

If the sin 90° is 1, then the cos ____ °= _____. 90° ; 1 over 2 0° ; square root 2 over 2 90° ; square root 3 over 2 0° ; 1

OpenStudy (alexatiger):

I'm sure the answer is the last one, but I don't know how to explain how i got the answer (especially because I'm guessing XD)

OpenStudy (anthonyym):

Yeah it is the last one

OpenStudy (alexatiger):

Can you explain to me how??

OpenStudy (anthonyym):

 1^2 = 0^2 + x^2 x = 1 So, cos(0) = 1